# Split Federated Learning for UAV-Enabled Integrated Sensing, Computation, and Communication

### v1

@@ V-A Experiment Settings
We consider a UAV-assisted ISCC system with $M$=5 UAVs. We set the sensing transmit power $p_{s}$=30 dBm, communication transmit power $p_{c}$=20 dBm, communication bandwidth $B$=1.2 MHz, noise power spectral density $N_{0}$=-174 dBm/Hz, height $H$=10 m, chirp duration $T_{d}$=10 $\mu$s, chirp numbers $\lambda$=25, environmental parameters $\varphi$=12.08 and $\varsigma$=0.1139. Then, we use the simulator proposed in[24] to generate five different human motion datasets: child/adult standing, child pacing, adult pacing, child walking, and adult walking. The spectrograms corresponding to different human motions can be seen in Fig. 2. We employ ResNet18 as the classification model. The network commences with a large-kernel convolution and downsampling, followed by four groups of residual blocks that utilize pre-activation units to enhance information flow and training efficiency. As the network progresses, it incrementally increases the number of channels while decreasing spatial resolution. High-level features are extracted through a global average pooling layer, and the predicted class is ultimately output by a fully connected layer.
To demonstrate the effectiveness of our proposed SFLSCC, we compared our scheme with the following schemes:
  • Ours: This scheme is based on the proposed SFLSCC framework and the control algorithm presented in Section IV, jointly optimizing the UAV positions $\left\{\boldsymbol{u}_{m}\right\}$, client aggregation frequency $I$, the data sensing volume $b$, and the split point $L_{c}$.
@@ V-B Experiment Analysis
To demonstrate the performance improvement of our proposed optimization algorithm, we compare the convergence and the required energy consumption to achieve the specified model accuracy among different schemes, as shown in Fig. 3. It shows that the proposed optimization algorithm significantly outperforms the baselines in terms of total energy consumption required for training, while achieving comparable convergence speed to baseline 3, which has the highest sensing quality.
As shown in Figs. 4 and 5, we respectively analyze the impact of different decision variables on the energy consumption and convergence of SFLSCC. As shown in Figs. 4(a) and  5(a), the higher the client-side aggregation frequency (i.e., the smaller the value of $I$), the faster the model converges and the lower the total energy consumption required to reach the specified model convergence accuracy. Such improvements are attributed to the fact that the differences among the client-side sub-models are diminished through more frequent aggregations. As shown in Fig. 5(b), the shallower the split point (i.e., the smaller the value of $L_{c}$), the faster the model converges. This is because the majority of the model is held by the edge server, which aggregates the model updates in each round. Compared to deeper split points, this results in a more stable convergence process. Additionally, a shallower split point means that UAVs possess fewer model layers, thereby reducing the computational burden and the communication overhead required to send the client-side model to the server for aggregation. This phenomenon is also confirmed in Fig. 4(b). As shown in Figs. 4(c) and 5(c), the larger the data sensing volume $b$, the faster the model converges, at the expense of increased energy consumption by the UAVs.
@@
The aforementioned experiments are all conducted under relatively ideal conditions. To investigate the impact of different environments on the performance of SFLSCC and thereby guide subsequent designs, we conduct experiments in suburban, urban, dense urban, and high-rise urban environments, as shown in Fig. 6. Complex environments can deteriorate model convergence, as occlusions in such environments affect the probability of Los links, thereby impacting the quality of UAV sensing and communication channels.
@@ VI Conclusion
This paper presents a novel SFL framework for UAV-enabled integrated sensing, computation, and communication, named SFLSCC, addressing key challenges in UAV-assisted FEL systems. By leveraging the unique capabilities of UAVs for local data collection and computation, SFLSCC optimizes the trade-off between training accuracy and energy consumption. Our theoretical analysis provides insights into the convergence behavior of SFLSCC, establishing upper bounds for the convergence gap based on UAV deployment, split point selection, and data sensing volume. We also propose a joint optimization approach to minimize energy consumption while ensuring the desired model accuracy. The non-convex nature of the problem is tackled with a low-complexity algorithm that efficiently delivers near-optimal solutions. Extensive experiments on a target motion recognition task demonstrate the superiority of SFLSCC in terms of convergence performance and energy efficiency compared to existing methods.
@@
=$\overset{\left(a\right)}{=}\frac{\eta}{2}\mathbb{E}\left[\left\|\sum\limits_{m=1}^{M}\varphi_{m}\left(\nabla\mathcal{L}\left(\boldsymbol{w}^{t-1}\right)-\nabla\mathcal{L}_{m}\left(\boldsymbol{w}_{m}^{t-1}\right)\right)\right\|^{2}\right.$
@@
$X_{B}=\frac{\eta}{2}\mathbb{E}\left[\left\|\sum\limits_{m=1}^{M}\varphi_{m}\left(\nabla\mathcal{L}\left(\boldsymbol{w}^{t-1}\right)-\nabla\mathcal{L}_{m}\left(\boldsymbol{w}^{t-1}\right)\right.\right.\right.$ (57)
@@
+$\varphi_{m}\left(\nabla\mathcal{L}_{m}\left(\boldsymbol{w}^{t-1}\right)-\nabla\mathcal{L}_{m}\left(\boldsymbol{w}^{t-1}_{m}\right)\right)\vphantom{\sum\limits_{m=1}^{M}}\right\|^{2}\right]$
@@
$\quad=\eta^{2}\mathbb{E}\left[\left\|\frac{\sum\limits_{M=1}^{m}\alpha_{m}^{t-1}\left(\mathbf{g}_{m}^{t-1}-\nabla\mathcal{L}_{m}(\boldsymbol{w}_{m}^{t-1})\right)}{M}\right.\right.$
@@ VI-1 Bound of $X$
$X$ =$\overset{\left(a\right)}{=}\mathbb{E}\left[\frac{\sum\limits_{m=1}^{M}\alpha_{m}^{t-1}\left\|\left(\mathbf{g}_{m}^{t-1}-\nabla\mathcal{L}_{m}(\boldsymbol{w}_{m}^{t-1})\right)\right\|^{2}}{M^{2}}\right]$ (60)
@@
=$\overset{\left(b\right)}{=}\frac{1}{M^{2}}\sum\limits_{m=1}^{M}\mathbb{E}\left[\left\|\left(\mathbf{g}_{m}^{t-1}-\nabla\mathcal{L}_{m}(\boldsymbol{w}_{m}^{t-1})\right)\right\|^{2}\right]$
@@ VI-2 Bound of $Y$
=$\overset{\left(b\right)}{=}2\underbrace{\mathbb{E}\left[\frac{\sum\limits_{m=1}^{M}\alpha_{m}^{t-1}\left\|\left(\nabla\mathcal{L}_{m}(\boldsymbol{w}^{t-1}_{m})-\nabla\mathcal{L}(\boldsymbol{w}^{t-1}_{m})\right)\right\|}{M^{2}}^{2}\right]}_{Y_{1}}$
@@
where $\left(a\right)$ follows from $\|\boldsymbol{x}+\boldsymbol{y}\|^{2}\leq 2\|\boldsymbol{x}\|^{2}+2\|\boldsymbol{y}\|^{2}$, equality $\left(b\right)$ is similar to Eq. 60$\left(a\right)$ and $\Phi_{i}=\nabla\mathcal{L}_{i}(\boldsymbol{w}^{t-1}_{i})-\nabla\mathcal{L}(\boldsymbol{w}^{t-1}_{i})$. Next, we find the upper bound of $Y_{1}$, $Y_{2}$ and $Y_{3}$ as follows.
@@
<$\overset{\left(a\right)}{<}\sum_{m=1}^{M}\kappa_{m}\left((q_{s,m}-\bar{q}_{s})^{2}+\bar{q}_{s}^{2}\right)\Lambda_{m}^{2},$
@@
$\quad+16\beta^{2}\eta^{2}I^{2}\sum\limits_{l=1}^{L_{c}}G_{l}^{2}+4\mathbb{E}\left[\|\nabla\mathcal{L}(\boldsymbol{w}^{t-1})\|^{2}\right].$
@@
$+32\beta^{2}\eta^{4}I^{2}\sum\limits_{l=1}^{L_{c}}G_{l}^{2}+8\eta^{2}\mathbb{E}\left[\|\nabla\mathcal{L}(\boldsymbol{w}^{t-1})\|^{2}\right].$
@@
$\quad+\frac{2\beta\eta^{2}}{M^{2}}\sum\limits_{m=1}^{M}\Lambda_{m}^{2}+16\beta^{3}\eta^{4}I^{2}\sum\limits_{l=1}^{L_{c}}G_{l}^{2}$
@@
$\quad+\frac{2\beta\eta^{2}}{M^{2}}\sum\limits_{m=1}^{M}\Lambda_{m}^{2}+16\beta^{3}\eta^{4}I^{2}\sum\limits_{l=1}^{L_{c}}G_{l}^{2}$
@@
$\quad+\frac{2\beta\eta^{2}}{M^{2}}\sum\limits_{m=1}^{M}\Lambda_{m}^{2}+16\beta^{3}\eta^{4}I^{2}\sum\limits_{l=1}^{L_{c}}G_{l}^{2}$

### v2

@@ V-A Experiment Settings
[…]
To demonstrate the effectiveness of our proposed SFLSC3, we compared our scheme with the following schemes:
  • Ours: This scheme is based on the proposed SFLSC3 framework and the control algorithm presented in Section IV, jointly optimizing the UAV positions $\left\{\boldsymbol{u}_{m}\right\}$, client aggregation frequency $I$, the data sensing volume $b$, and the split point $L_{c}$.
@@
  • Baseline 4: This scheme considers that the UAV positions $\left\{\boldsymbol{u}_{m}\right\}$ are fixed, and jointly optimizes the client aggregation frequency $I$, the data sensing volume $b$, and the split point $L_{c}$, as discussed in Section IV.
@@ V-B Experiment Analysis
To demonstrate the performance improvement of our proposed optimization algorithm, we compare the convergence and the required delay to achieve the specified model accuracy among different schemes, as shown in Fig. 3. It shows that the proposed optimization algorithm significantly outperforms the baselines in terms of total training delay, while achieving comparable convergence speed to baseline 3, which has the highest sensing quality.
As shown in Figs. 4 and 5, we investigate the effects of key decision variables on both the convergence behavior and delay performance of SFLSC3. The results indicate that none of the control factors exhibits a purely monotonic impact on system performance; instead, clear trade-offs can be observed. First, as illustrated in Figs. 4(a)5(a) and 6(a), increasing the client-side aggregation frequency (i.e., decreasing $I$) indeed accelerates model convergence in terms of the number of communication rounds. This improvement can be attributed to the fact that more frequent aggregations effectively reduce the divergence among client-side sub-models. However, each aggregation operation incurs additional communication and computation overhead, leading to a higher average per-round delay. Consequently, when evaluated from a long-term perspective, a smaller $I$ does not necessarily result in a shorter wall-clock time to reach a target accuracy. Instead, an intermediate aggregation interval achieves a more favorable trade-off between convergence speed and overall delay. Second, the impact of the split point $L_{c}$ exhibits a similar non-monotonic behavior. As shown in Figs. 4(b),  5(b) and 4(b), a shallower split point (i.e., a smaller $L_{c}$) leads to faster convergence in terms of rounds, since the majority of the model parameters are maintained and aggregated at the edge server, resulting in more stable updates. Nevertheless, the delay performance does not monotonically improve as $L_{c}$ decreases. When $L_{c}$ approaches the boundary values, the delay performance deteriorates. This is because although a smaller $L_{c}$ reduces the size of the client-side model to be transmitted, it simultaneously increases the volume of intermediate feature data that must be exchanged at the split layer, which may dominate the communication overhead. These observations indicate that the effects of $I$ and $L_{c}$ on convergence and delay are inherently coupled. Third, as shown in Figs. 4(c),  5(c) and 6(c), increasing the data sensing volume $b$ accelerates convergence in terms of rounds, owing to the improved data diversity and learning effectiveness at each UAV. However, this comes at the cost of increased sensing, processing, and transmission delay at the UAV side. Interestingly, under the considered experimental settings, a smaller $b$ yields better delay performance, which is mainly due to the relatively simple learning task. Nevertheless, when $b$ becomes excessively small, the convergence performance degrades significantly, revealing another clear trade-off between learning efficiency and delay. Finally, we examine the influence of UAV positions $\left\{\boldsymbol{u}_{m}\right\}$, characterized by the sensing angle $\theta_{s}$. Figs. 4(d),  5(d) and 6(d) show that increasing $\theta_{s}$ improves the sensing success probability, but also leads to higher delay due to the enlarged sensing and communication overhead. Therefore, $\theta_{s}$ is not always preferable when set to larger values, and an appropriate balance must be achieved. In summary, all the considered control variables, including $I$, $L_{c}$, $b$, and $\theta_{s}$, involve inherent trade-offs between convergence efficiency and delay performance. These observations further demonstrate the necessity and practicality of the proposed optimization framework.
@@
The aforementioned experiments are all conducted under relatively ideal conditions. To investigate the impact of different environments on the performance of SFLSC3 and thereby guide subsequent designs, we conduct experiments in suburban, urban, dense urban, and high-rise urban environments, as shown in Fig. 7. Complex environments can deteriorate model convergence, as occlusions in such environments affect the probability of Los links, thereby impacting the quality of UAV sensing and communication channels.
@@ VI Conclusion
This paper presents a novel SFL framework for UAV-enabled integrated sensing, computation, and communication, named SFLSC3, addressing key challenges in UAV-assisted FEL systems. By leveraging the unique capabilities of UAVs for local data collection and computation, SFLSC3 optimizes the trade-off between training accuracy and delay. Our theoretical analysis provides insights into the convergence behavior of SFLSC3, establishing upper bounds for the convergence gap based on UAV deployment, split point selection, and data sensing volume. We also propose a joint optimization approach to minimize delay while ensuring the desired model accuracy. The non-convex nature of the problem is tackled with a low-complexity algorithm that efficiently delivers near-optimal solutions. Extensive experiments on a target motion recognition task demonstrate the superiority of SFLSC3 in terms of convergence and latency performance compared to existing methods.
@@
$\overset{\left(a\right)}{=}\frac{\eta}{2}\mathbb{E}\left[\left\|\sum\limits_{m=1}^{M}\varphi_{m}\left(\nabla\mathcal{L}\left(\boldsymbol{w}^{t-1}\right)-\nabla\mathcal{L}_{m}\left(\boldsymbol{w}_{m}^{t-1}\right)\right)\right\|^{2}\right$.
@@
[…]
@@
[…]
@@
[…]
@@ VI-1 Bound of $X$
$X$ $\overset{\left(a\right)}{=}\mathbb{E}\left[\frac{\sum\limits_{m=1}^{M}\alpha_{m}^{t-1}\left\|\left(\mathbf{g}_{m}^{t-1}-\nabla\mathcal{L}_{m}(\boldsymbol{w}_{m}^{t-1})\right)\right\|^{2}}{M^{2}}\right]$ (56)
@@
$\overset{\left(b\right)}{=}\frac{1}{M^{2}}\sum\limits_{m=1}^{M}\mathbb{E}\left[\left\|\left(\mathbf{g}_{m}^{t-1}-\nabla\mathcal{L}_{m}(\boldsymbol{w}_{m}^{t-1})\right)\right\|^{2}\right]$
@@ VI-2 Bound of $Y$
$\overset{\left(b\right)}{=}2\underbrace{\mathbb{E}\left[\frac{\sum\limits_{m=1}^{M}\alpha_{m}^{t-1}\left\|\left(\nabla\mathcal{L}_{m}(\boldsymbol{w}^{t-1}_{m})-\nabla\mathcal{L}(\boldsymbol{w}^{t-1}_{m})\right)\right\|}{M^{2}}^{2}\right]}_{Y_{1}}$
@@
[…]
@@
$\overset{\left(a\right)}{<}\sum_{m=1}^{M}\kappa_{m}\left((q_{s,m}-\bar{q}_{s})^{2}+\bar{q}_{s}^{2}\right)\Lambda_{m}^{2},$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]